### v1

# The MASSIVE Survey. XX. A Triaxial Stellar Dynamical Measurement of the Supermassive Black Hole Mass and Intrinsic Galaxy Shape of Giant Radio Galaxy NGC 315

Abstract
[…]
galaxies: elliptical and lenticular, cD — galaxies: evolution — galaxies: kinematics and dynamics — galaxies: stellar content — galaxies: structure — dark matter
software: Astropy (Astropy Collaboration et al., 2013, 2018), Dynesty (Speagle, 2020), Galfit (Peng et al., 2002), jampy (Cappellari, 2008), linmix (Kelly, 2007), Matplotlib (Hunter, 2007), mgefit (Cappellari, 2002), NumPy (Harris et al., 2020).
1 Introduction
The MASSIVE survey is a volume-limited, photometric and spectroscopic survey of the ${\sim}100$ most massive early-type galaxies (with stellar mass $M_{*}\gtrsim 10^{11.5}M_{\odot}$) in the local Universe (Ma et al., 2014). A key scientific goal of the survey is to dynamically measure the masses of a sample of supermassive black holes (SMBHs) within the targeted volume (to a distance of ${\sim}100$ Mpc above declination $\delta=-6^{\circ}$) with spatially resolved stellar and gas (when present) kinematics. To date, 14 MASSIVE galaxies have published dynamical SMBH mass ($M_{\mathrm{BH}}$) measurements as compiled in Table 1 of Liepold & Ma (2024). Among them, only M87 has $M_{\mathrm{BH}}$ determined from the motion of more than one type of dynamical tracers. For the rest, nine galaxies111 NGC 708, NGC 1453, NGC 1600, NGC 2693, NGC 3842, NGC 4472, NGC 4649, NGC 4889, NGC 7619 have $M_{\mathrm{BH}}$ determined from stellar kinematics using the Schwarzschild orbit modeling method, three galaxies222 NGC 315, NGC 997, NGC 1684 have $M_{\mathrm{BH}}$ inferred from CO kinematics, and one galaxy [^3] NGC 7052 is studied with ionized gas.
In this work, we report a new mass measurement of the SMBH in NGC 315 using stellar kinematics from MASSIVE survey observations and the triaxial orbit modeling method. NGC 315 has a prior $M_{\mathrm{BH}}$ determination based on CO kinematics from Atacama Large Millimeter/submillimeter Array (ALMA) observations (Boizelle et al., 2021). NGC 315 is therefore only the second galaxy in the MASSIVE survey for which a direct comparison of $M_{\mathrm{BH}}$ from different dynamical tracers can be made. Beyond MASSIVE, six other galaxies have $M_{\mathrm{BH}}$ inferred from both ALMA CO kinematics and stellar kinematics, enabling us to assess the consistency between the two methods (see Sec. 4). NGC 315 is also only the 5th MASSIVE galaxy for which the triaxial stellar orbit modeling is used to determine its $M_{\mathrm{BH}}$ (others assumed axisymmetry), and the first MASSIVE galaxy with $M_{\mathrm{BH}}$ determined from both CO and triaxial stellar based methods.
NGC 315 is the brightest member of a galaxy group identified in the Two Micron All Sky Survey (2MASS) group catalog (Crook et al., 2007). The high-density contrast and low-density contrast versions of the catalog list 6 and 97 member galaxies, respectively. The halo virial mass is estimated to be $3.5{\times}10^{13}M_{\odot}$ based on member galaxy velocities. NGC 315 has strong nuclear radio emission (Fanaroff & Riley, 1974) and a prominent jet extending $\gtrsim 100^{\prime\prime}$ at a position angle (PA) of ${\sim}-50^{\circ}$ (Laing et al., 2006; Ricci et al., 2022). NGC 315 is one of seven MASSIVE survey targets with evidence for an X-ray point source in the nuclear region in the $4-7$ keV band. The mean temperature and luminosity of the X-ray hot gas are estimated to be $T_{X}=0.57$ keV and $L_{X}=3.8{\times}10^{41}$ erg s-1 (Goulding et al., 2016).
Typical of MASSIVE galaxies, NGC 315 is a slow rotator with a velocity amplitude of ${\sim}30{\rm\,km~{}s^{-1}}$ and a spin parameter of $\lambda=0.063$ (Veale et al., 2017b; Ene et al., 2019). The PA of the kinematic axis (measured E of N to the receding portion) is determined to be $\text{PA}_{\text{kin}}=222^{\circ}{\pm}7^{\circ}$ over the $107^{\prime\prime}{\times}107^{\prime\prime}$ field of view (FOV) of the Mitchell integral-field spectrograph (IFS, Ene et al. 2018), and $\text{PA}_{\text{kin}}=218^{\circ}{\pm}13^{\circ}$ in the central $5^{\prime\prime}{\times}7^{\prime\prime}$ region from Gemini IFS data (Ene et al., 2020). The ${\sim}90^{\circ}$ offset between the kinematic axis and the jet indicates the projection of the angular momentum vector of the stars onto the sky is at the same PA as the jet. HST photometry shows boxy isophotes and nearly constant ellipticity and photometric PA between a radius of $1^{\prime\prime}$ and $100^{\prime\prime}$ with luminosity-weighted values of $\epsilon=0.27$ and $\text{PA}_{\text{phot}}=44.3^{\circ}{\pm}0.2^{\circ}$ (Fig. 1.2 of Goullaud et al. 2018). The kinematic misalignment angle is consistent with 0: $\Psi=6.3^{\circ}{\pm}13.3^{\circ}$ (Ene et al., 2020).
In Section 2, we discuss the IFS data from Gemini and McDonald Observatories and the stellar velocity moment measurements reported in Veale et al. (2017b, a, 2018); Ene et al. (2018, 2019, 2020). The HST observations of NGC 315 (Goullaud et al., 2018) and surface brightness profile determination (Boizelle et al., 2021) are also summarized. In Section 3, we summarize the triaxial orbit modeling code TriOS (Quenneville et al., 2021, 2022) and the parameter search strategy used to select galaxy models, followed by a discussion of the mass and shape parameters and stellar orbital structure that best match the observations. Section 4 discusses systematic uncertainties and compares $M_{\mathrm{BH}}$ determined from stellar vs. CO kinematics.
Throughout this work, we assume a luminosity distance for NGC 315 of $D_{L}=68.1{\pm}2.5$ Mpc from the MASSIVE-HST project using the surface brightness fluctuation technique (Goullaud et al., 2018; Jensen et al., 2021; Blakeslee et al., 2021). At NGC 315’s redshift of $z=0.0165$, the corresponding angular diameter distance is $D_{A}=65.9\pm 2.4$ Mpc, and $1^{\prime\prime}$ is $320$ pc. When comparing with Boizelle et al. (2021), we adjust their reported values from their assumed $D_{A}=70$ Mpc to our distance.
2 Observations
2.1 Photometry
We adopt the characterization of the surface brightness of NGC 315 by Boizelle et al. (2021) based on archival HST and Spitzer photometry. The HST Wide Field Camera 3 (WFC3) observation in the F110W filter cover a $2.1^{\prime}{\times}2.2^{\prime}$ region centered at NGC 315 and produce a final image with $0\farcs 08\text{ pixel}^{-1}$ scale (GO-14219; Goullaud et al. 2018). Archival Spitzer InfraRed Array Camera data from channel 1 ($3.6\mu m)$ provide deeper coverage of the stellar halo. The final mosaiced image covers a radial range out to $R{\sim}11^{\prime}$ (about $225$ kpc).
NGC 315 has a prominent circumnuclear dust disk in the central ${\sim}1^{\prime\prime}$ region. Boizelle et al. (2021) produce three multi-Gaussian expansion (MGE) models (A, B2, and B3) of the mosaiced image, applying a differing amount of extinction correction to the nuclear region. The central Gaussian component in B2 and B3 has a width of $\sigma^{\prime}=0\farcs 178$ and $\sigma^{\prime}=0\farcs 119$, respectively, comparable to the PSF (${\sim}0\farcs 15$). As discussed in Liepold et al. (2025), the width of the central component is poorly constrained when it is comparable or below the scale of the PSF. The MGE routine sometimes assigns a central component with $\sigma^{\prime}\lesssim\sigma^{\prime}_{\rm PSF}$ that does not improve the fit in a meaningful way but results in an exceptionally large central 3D density after deprojection (recall $\nu\propto\Sigma/\sigma^{\prime}$ where $\Sigma$ is the central surface brightnes). Accordingly, the B2 and B3 MGEs have a “bump” in their luminosity densities in the central region, as shown in Figure 7 in the Appendix. Additional examples of this artifact can be found in Davidson et al. (2024). Liepold et al. (2025) circumvented this problem by imposing a lower limit on $\sigma^{\prime}$ during the MGE fitting. Here, we adopt MGE model A, which has $\sigma^{\prime}=0\farcs 580$ for the central component and a smooth 3D deprojected density profile without this stellar excess. In Sec. 4.1 below, we describe results from tests using model B3, the case with the most extreme central luminosity density.
Figure 1: Four representative Gemini GMOS spectra (black) of NGC 315 for spatial bins located at increasing distance from the nucleus. The stellar template broadened by the best-fit LOSVD is overlaid (blue) on each spectrum. The fitting residuals (red points) are offset by constants for clarity. The typical residual is ${\sim}0.5\%$. The grey shaded regions are excluded from the fit to account for improperly subtracted sky lines and detector gap.
2.2 Integral-Field Spectroscopy
Figure 2: Stellar kinematic maps of the central $5^{\prime\prime}\times 7^{\prime\prime}$ region of NGC 315 from Gemini GMOS observations. Spectra from individual lenslets are co-added to achieve a single spectrum with $S/N\gtrsim 125$ for each of the 245 spatial bins. The two upper-left panels show the line-of-sight velocities $V$ and velocity dispersions $\sigma$, with the higher-order Gauss-Hermite moments $h_{3}$ to $h_{8}$ shown in the other panels. Surface brightness contours are plotted as dotted gray lines.
2.2.1 Central stellar kinematics
We observed the central $5^{\prime\prime}{\times}7^{\prime\prime}$ of NGC 315 using the two-slit IFS mode of Gemini Multi-Object Spectrograph (GMOS) on the Gemini North Telescope with 1000 hexagonal lenslets (each with a projected diameter of $0\farcs 2$). Ten science exposures of 1200 seconds each were obtained, totaling 3.3 hours of on-source and simultaneous observations of a $5^{\prime\prime}{\times}3\farcs 5$ region of sky offset by ${\sim}1^{\prime}$ from the galaxy. The R400400400400-G5305530553055305 grating with the CaT filter was used to cover the wavelength range $7800{-}9330$. The median seeing was $0\farcs 6$ FWHM, and the median spectral resolution (determined from arc lamp lines for each lenslet) was $2.5$ FWHM. Details of the data processing procedure are described in MASSIVE Paper XIII (Ene et al., 2019).
We co-add the spectra from a group of adjacent GMOS lenslets to achieve a threshold signal-to-noise ratio ($S/N$) of 125. This binning procedure results in 248 high-quality spectra covering the central region of NGC 315. One difference in this step from Ene et al. (2019) is that a symmetric binning scheme over the four quadrants of the galaxy was used in that work, while here we perform the binning over the entire GMOS FOV without this assumption. This difference only introduces minor adjustments in how the GMOS lenslets are grouped spatially; the stellar kinematics are measured for each bin without any symmetry assumption in both analyses.
From each spectrum, we measure the line-of-sight velocity distribution (LOSVD) from the CaII triplet (CaT) absorption features over a rest wavelength range of $8420{-}8770$ using the penalized pixel-fitting (pPXF) method (Cappellari, 2017). The LOSVD is represented as a Gauss-Hermite series of order $n=8$. We use $15$ stellar templates from the MILES CaT Library that covers a wavelength range of $8437{-}9020$ with a spectral resolution of $1.5$ FWHM (Cenarro et al., 2001). This set of stars is taken from Table 2 of Barth et al. (2002), but we find consistent stellar kinematics when the full library of $706$ stars is used. A multiplicative polynomial of degree three is used to model the stellar continuum in each spectrum.
Four representative spectra at increasing radii are plotted in Figure 1. The template spectrum broadened by the best-fit LOSVD (blue curve) provides an excellent fit to each observed spectrum (black curve) with a typical residual (red points) of ${\sim}0.5$%. In addition to the bright sky lines, the wavelength range ${\gtrsim}8750$ is also masked due to the presence of a gap in the CCD chip. The resulting maps of the eight Gauss-Hermite moments of the LOSVDs for the 248 GMOS spatial bins are shown in Figure 2. The corresponding radial profiles of the eight moments are displayed as blue bars in Figure 3. The errors on the moments are determined via the Monte Carlo method described in Section 4 of Ene et al. (2019). These figures show a velocity profile with a low amplitude rotational velocity $|V|{\sim}20\,{\rm\,km~{}s^{-1}}$, and a velocity dispersion profile that rises from $\sigma{\sim}320{\rm\,km~{}s^{-1}}$ at $R{\sim}2^{\prime\prime}$ inward to $\sigma{\sim}350{\rm\,km~{}s^{-1}}$ at $R{\sim}0\farcs 3$.
Figure 3: Radial profiles of the stellar kinematic moments for NGC 315 from Gemini GMOS (blue bars) and Mitchell (pink bars) data. Moments predicted by the best-fitting triaxial galaxy model (black squares) listed in Table 1 match the input data well.
2.2.2 Wide-field stellar kinematics
[…]
The LOSVD from each Mitchell spectrum is extracted in a similar way as GMOS described above. We opt to fit to $n=6$ Gauss-Hermite moments due to the lower $S/N$ data here. The MILES library of $985$ stellar spectra is used as templates (Sanchez-Blazquez et al., 2006; Falcón-Barroso et al., 2011). The kinematic moments for the 55 Mitchell bins are shown in Figure 3 in pink. The Mitchell data points connect smoothly to the GMOS points, showing excellent agreement between measurements obtained from different spectrographs, telescopes, and spectral regions. Six additional kinematic points at $R{\sim}50^{\prime\prime}$ are shown in Figure D1 of Veale et al. (2017b) but are excluded in the following analysis due to the low $S/N$ of these outer spectra.
3 Results from Triaxial Orbit Modeling
3.1 The TriOS Code and Galaxy Models
We use the TriOS code (Quenneville et al., 2021, 2022) to compute triaxial orbit models of NGC 315. This code integrates a large number of stellar orbits that span the allowed phase space and computes the LOSVDs for a wide range of galaxy model parameters. The galaxy is assumed to have three mass components: a central black hole of mass $M_{\mathrm{BH}}$, a stellar component with a mass-to-light ratio $M^{*}/L$, and a dark matter halo with a density profile $\rho(r)=\rho_{0}/[(r/r_{s})^{\gamma}\left(1+r/r_{s}\right)^{3-\gamma}]$, where $r_{s}$ is a scale radius (Navarro et al., 1996). We set $\gamma=0$ so that the profile has a finite central density $\rho_{0}$ and a flattened central density distribution similar to that of the stars. With our data, $r_{s}$ and $\rho_{0}$ are often quite degenerate so we choose to parameterize the halo with a single parameter, $M_{15}$, defined to be the dark matter mass enclosed within $15$ kpc with a fixed scale radius $r_{s}=15$ kpc. A similar strategy was used in orbit modeling of other MASSIVE galaxies (e.g., Liepold et al. 2020; Quenneville et al. 2022; Pilawa et al. 2022).
We use three parameters to specify the triaxial shape of the stellar component: $p=b/a$ is the intrinsic middle-to-long axis ratio, $q=c/a$ is the intrinsic short-to-long axis ratio, and $u$ is the apparent-to-intrinsic long axis ratio. These three shape parameters are related to the three angles $\theta$, $\phi$, and $\psi$ that relate the intrinsic and projected coordinate systems of NGC 315; see Equations (4) and (8) of Quenneville et al. (2022). Here $\theta$ and $\phi$ are polar angles in NGC 315’s intrinsic coordinate system, and $\psi$ specifies the remaining degree of freedom – a rotation of the galaxy around the line of sight.
[…]
Figure 4: (Left) Posterior distributions of six parameters in triaxial orbit modeling of NGC 315: SMBH mass $M_{\mathrm{BH}}$, stellar mass-to-light ratio $M^{*}/L$, dark matter enclosed within 15 kpc $M_{15}$, and luminosity averaged axis ratios $u,p$, and $q$. The 68%, 95%, and 99.7% credible regions are represented by different shades of purple. The vertical lines in each 1D marginalized distribution indicate the median and the three corresponding confidence levels. (Upper right) Posterior distributions in viewing angle space, where $\theta$ and $\phi$ are polar angles in the galaxy’s frame, and $\psi$ specifies a rotation of the galaxy around the line of sight (Section 2.2 of Quenneville et al. 2022).
Table 1
Mass and Shape Parameters of NGC 315
Galaxy Property (units) Value
Black hole mass $M_{\mathrm{BH}}$ $(10^{9}M_{\odot})$ $3.0{\pm}0.3$
$M^{*}/L$ ($M_{\odot}/L_{\odot}$) $2.49{\pm}0.06$
DM mass within 15 kpc $M_{15}$ $(10^{11}M_{\odot})$ $9.9^{+0.9}_{-0.7}$
Triaxiality parameter $T$ $0.89{\pm}0.02$
Shape parameter $T_{\text{maj}}$ $0.019^{+0.008}_{-0.007}$
Shape parameter $T_{\text{min}}$ $0.017^{+0.009}_{-0.013}$
Average middle-to-long axis ratio $p$ $0.854^{+0.004}_{-0.003}$
Average short-to-long axis ratio $q$ $0.833^{+0.002}_{-0.001}$
Average apparent-to-intrinsic long axis ratio $u$ $0.998{\pm}0.001$
Line-of-sight direction $\theta(^{\circ})$, $\phi(^{\circ})$ $82.6^{+2.3}_{-2.5}$, $82.0^{+1.8}_{-1.4}$
Rotation about line of sight $\psi\left({}^{\circ}\right)$ $-89.1^{+0.4}_{-0.3}$
Table 1: For each parameter, we marginalize over the other parameters and report the 68% credible regions. In orbit models, $\theta$ is the inclination angle in the oblate axisymmetric limit ($\psi=90^{\circ}$, or equivalently $p=1$), with $\theta=90^{\circ}$ being edge-on and $\theta=0^{\circ}$ being face-on.
3.2 Best-Fit Triaxial Model
We use the model selection scheme described in our recent work (e.g., Pilawa et al. 2022; Liepold et al. 2023, 2025) to determine the mass and shape parameters that best match the kinematic and photometric data of NGC 315. This scheme, in brief, involves generating candidate galaxy models with Latin hypercube sampling (McKay et al., 1979), approximating the resulting likelihood landscapes with Gaussian process regression, and sampling posteriors on our parameters via dynamic nested sampling (Speagle, 2020). About ${\sim}1750$ galaxy models are used to obtain the final posteriors on the six parameters shown in Figure 4. An additional ${\sim}3000$ models covering wider ranges of parameters are used in the initial exploration of the 6D likelihood surface, and ${\sim}10,000$ more models are used in various tests to ensure the surface in the low $\chi^{2}$ region is mapped out accurately.
A summary of the best-fit parameters for NGC 315 is listed in Table 1. The kinematic moments predicted by the model are compared with observed values in Figure 3. The total $\chi^{2}$ is $1611.4$ spread over $2314$ kinematic constraints ($8$ moments for 248 GMOS bins and $6$ moments for 55 Mitchell bins). A naive estimate of the reduced $\chi^{2}$ would be $1611.4/(2314-6)=0.698$ assuming 6 degrees of freedom (DOF). But we caution that DOF is nontrivial to estimate for nonlinear problems such as here. In a study using simulated stellar kinematics that mimic data in the MASSIVE survey (Pilawa et al., 2024), we have investigated a “generalized” measure of DOF for 6-parameter models and found it to be ${\sim 200}$ instead of the canonical value of 6, thereby raising the reduced $\chi^{2}$ by ${\sim}40$% in that study. While a similar calculation would have to be performed to estimate this DOF measure for NGC 315, we expect the reduced $\chi^{2}$ to be raised as well.
3.2.1 Black Hole Mass
To assess how NGC 315 and its SMBH fit in the population of local galaxies with dynamically inferred $M_{\mathrm{BH}}$, we examine its location on the well-studied $M_{\mathrm{BH}}{-}\sigma$ and $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations. NGC 315’s luminosity-weighted velocity dispersion within $R_{e}$ is found to be $\sigma_{e}=341\text{ km s}^{-1}$ based on the same Mitchell IFS data in this paper (Veale et al., 2017b). At this $\sigma_{e}$, the mean $M_{\mathrm{BH}}{-}\sigma$ relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=4.2\times 10^{9}M_{\odot}$ and $3.3\times 10^{9}M_{\odot}$, respectively, $40\%$ (0.15 dex) and $10\%$ (0.04 dex) larger than our dynamically measured $M_{\mathrm{BH}}$. However, our $M_{\mathrm{BH}}$ is within the intrinsic scatter of both relations, $0.38$ dex.
To place NGC 315 on the $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relation, we use the total stellar mass from our best-fitting triaxial model, $M_{*}=1.5\times 10^{12}M_{\odot}$. At this bulge mass, the mean relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=5.3\times 10^{9}M_{\odot}$ and $4.4\times 10^{9}M_{\odot}$, respectively, $77$% (0.25 dex) and $47$% (0.17 dex) larger than our measured $M_{\mathrm{BH}}$. But again, our dynamical $M_{\mathrm{BH}}$ is within the intrinsic scatter of both $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations, $0.34$ dex.
3.2.2 Intrinsic 3D Galaxy Shape
The preferred axis ratios of $p=0.854$ and $q=0.833$ yield a triaxiality parameter of $T=\left(1-p^{2}\right)/\left(1-q^{2}\right)=0.89{\pm}0.02$, where the limits of $T=0$ and 1 correspond to oblate axisymmetry and prolate axisymmetry, respectively. Thus, NGC 315 is triaxial and highly prolate. In comparison, five other massive elliptical galaxies for which we have performed triaxial orbit modeling thus far all have smaller $T$. Four of them are oblate ($T\lesssim 0.5$): $T=0.33{\pm}0.06$ for NGC 1453 (Quenneville et al., 2021), $T=0.39{\pm}0.04$ for NGC 2693 (Pilawa et al., 2022), $T=0.35{\pm}0.03$ for Holmberg 15A (Liepold et al., 2025), and $T=0.31{\pm}0.05$ for NGC 57 (Pilawa et al. 2025, in prep). M87, on the other hand, is slightly prolate with $T=0.65{\pm}0.02$ (Liepold et al., 2023).
While the determinations of the intrinsic axis ratios $p$ and $q$ for individual galaxies require triaxial orbit modeling and detailed kinematic data, one can infer the distributions of $p$ and $q$ statistically from the observed isophotal shapes and kinematic vs. photometric misalignment angles of an ensemble of galaxies. For slow-rotating galaxies in the MASSIVE survey, the mean values are found to be $\langle p\rangle=0.88$, $\langle q\rangle=0.65$, and $\langle T\rangle=0.39$ (Ene et al., 2018), similar to the majority of our directly measured $T$ thus far. Comparable distributions are also found for early-type galaxies in the ATLAS3D and SAMI surveys (Weijmans et al., 2014; Foster et al., 2017) and for simulated massive slow rotators in the IllustrisTNG simulations (Pulsoni et al., 2020). We therefore expect stellar halos with the prolateness of NGC 315 to be a rare occurrence.
3.2.3 Stellar Orbital Structure
The composition of the major types of orbits in the best-fit model for NGC 315 is plotted as a function of radius in the upper panel of Figure 5. The two types of tube orbits have a fixed sense of rotation, with short-axis tubes having an angular momentum component along the minor axis that does not change sign, and similarly, long-axis tubes having an angular momentum component along the major axis that does not change sign. For the box orbits, all three components of the angular momentum change sign, leaving no sense of rotation for this orbit type. The relative contribution of these three orbit types determines the velocity structure of the galaxy.
Figure 5: Composition of the three major orbit types (upper panel) and velocity anisotropy (lower panel) as a function of radius in the best-fit triaxial galaxy model of NGC 315. The majority of the orbital weights are in long-axis tube orbits, consistent with the prolateness of the galaxy. Box orbits and short-axis tube orbits contribute about $25\%$ and $20\%$ of the remaining orbital weights, respectively. The velocity anisotropy parameters $\beta\equiv 1-\sigma_{t}^{2}/\sigma_{r}^{2}$ (in spherical coordinates) and $\beta_{z}=1-\sigma^{2}_{z}/\sigma^{2}_{R}$ (in cylindrical coordinates) indicate the stellar orbits are mildly tangentially anisotropic at small radii and are increasingly radially anisotropic at larger radii.
Figure 5 shows that beyond the central ${\sim}200\text{ pc}$, the fraction of long-axis tubes, short-axis tubes, and box orbits is nearly a constant of ${\sim}55\%$, ${\sim}20\%$, and ${\sim}25\%$, respectively. The predominance of long-axis tubes over short-axis tubes reflects the prolateness of NGC 315. This is opposite to the orbital compositions in oblate triaxial galaxies with $T<0.5$, e.g., about $60{-}70$% of orbits are short-axis tubes while ${\sim}20$% are long-axis tubes in NGC 1453 and NGC 2693 (Quenneville et al., 2021; Pilawa et al., 2022).
The lower panel of Figure 5 displays the radial profile of the velocity anisotropy parameters, $\beta=1-\sigma_{t}^{2}/\sigma_{r}^{2}$ and $\beta_{z}=1-\sigma_{z}^{2}/\sigma_{R}^{2}$, where $\sigma_{t}$ and $\sigma_{r}$ are the tangential and radial velocity dispersions in spherical coordinates, and $\sigma_{z}$ and $\sigma_{R}$ are the vertical and radial velocity dispersions in cylindrical coordinates, respectively. The orbits in the central portion of NGC 315 are slightly tangential with $\beta<0$ and become radially anisotropic away from the center. This trend is typical in massive elliptical galaxies (e.g., Thomas et al. 2016) and can be seen in other MASSIVE galaxies (e.g., Liepold et al. 2020; Pilawa et al. 2022).
4 Discussion
4.1 Uncertainties in Central Surface Brightness
A primary systematic uncertainty in the mass measurements of NGC 315 SMBH is the effect of dust on the observed central stellar light. In the main analysis above, we have used MGE model A of Boizelle et al. (2021) to approximate the surface brightness profile of NGC 315. To test the impact of the adopted profile on the inferred $M_{\mathrm{BH}}$, we replace it with their MGE B3, the model that assumes the largest extinction correction. We re-run orbit models and parameter search in the reduced 3D space of $M_{\mathrm{BH}}$, $M^{*}/L$, and halo mass with the shape parameters fixed to the best-fitting values. While model B3 has a significantly higher central luminosity density than model A (Figure 7), the difference is confined to the inner ${\sim}0\farcs 5$, a scale comparable to the PSF of our Gemini observations. It is thus not surprising that our tests find ${\lesssim 3}$% changes in the best-fitting parameters, with $M_{\mathrm{BH}}=(2.9\pm 0.2)\times 10^{9}M_{\odot}$ and $M^{*}/L=2.45\pm 0.04\,M_{\odot}/L_{\odot}$. In comparison, Boizelle et al. (2021) find that switching from model A to B3 in their CO-based study reduces $M_{\mathrm{BH}}$ and $M^{*}/L$ by ${\sim}18\%$. One possible reason for this larger change in their tests is the CO kinematic data only extend to ${\sim}1^{\prime\prime}$, and thus their constraints are more sensitive to changes in the central luminosity density and the enclosed mass.
Figure 6: Comparison of seven local galaxies with $M_{\mathrm{BH}}$ determined independently from stellar and CO kinematics listed in Table 2. The median linear fit (solid blue line) is given by $\log_{10}(M_{\mathrm{BH}}^{\rm CO}/10^{9}M_{\odot})=(1.15^{+0.44}_{-0.41})\log_{10}(M_{\mathrm{BH}}^{\rm stellar}/10^{9}M_{\odot})-(0.12{\pm}0.15)$, with an intrinsic scatter of $\epsilon=0.35^{+0.26}_{-0.12}\text{ dex}$ (dotted blue lines). Faint gray lines are constructed from sampling over the posteriors for our parameters. The solid black line denotes the one-to-one line to guide the eye. The two points for NGC 4751 represent $M_{\mathrm{BH}}$ inferred when assuming a spatially varying (lower point) vs. constant (upper point) $M^{*}/L$ in the CO-based model (see text).
4.2 $M_{\mathrm{BH}}$ from Stellar vs. CO Kinematics
With the mass measurement of NGC 315’s SMBH presented in this paper, there are now seven galaxies whose SMBH masses have been determined from the kinematics of both CO and stars (with orbit-based modeling). A summary of these measurements is given in Table LABEL:tab:summary and Figure 6. To assess the level of agreement between $M_{\mathrm{BH}}$ from the two methods, we perform a linear fit using the LinMix package (Kelly, 2007). This method derives the best-fitting line through hierarchical Bayesian modeling, accounting for uncertainties in both $x$ and $y$ directions. It models the distribution of the independent variable as a mixture of Gaussians and assumes the dependent variable is drawn from a Gaussian distribution centered on a linear relation with the independent variable. The code returns posterior distributions for the parameters, including an intrinsic scatter term representing the variance beyond measurement uncertainties.
The resulting linear fit for $M_{\mathrm{BH}}$ from the two methods is $\log_{10}(M_{\mathrm{BH}}^{\rm CO}/10^{9}M_{\odot})=(1.15^{+0.44}_{-0.41})\log_{10}(M_{\mathrm{BH}}^{\rm stellar}/10^{9}M_{\odot})-(0.12{\pm}0.15)$, with an intrinsic scatter of $\epsilon=0.35^{+0.26}_{-0.12}$ dex (blue lines in Figure 6). To assess whether the data prefer a one-to-one model with a fixed slope of $1$ and intercept of 00 or a linear relation with free coefficients, we compare the marginal evidence ratio (“Bayes factor”) of the two models. The one-to-one model here is a subset of the free linear model, so the Bayes factor in this case quantifies the factor by which a free slope and intercept improves the fit. Using dynesty, we find a log-marginal evidence difference of $\Delta\ln Z=2.78\pm 0.02$ (one-to-one model minus free model), which is much less than what is typically considered evidence in favor of one model (e.g., $|\Delta\ln Z|>5$, Trotta 2008; Lockhart & Gralla 2022). We therefore conclude that while there are differences in individual $M_{\mathrm{BH}}$ values determined from CO and stellar kinematics, there is no evidence for statistically significant biases between the two methods in the current data.
Table LABEL:tab:summary
Galaxies with dynamical $M_{\mathrm{BH}}$ measurements from both stellar and CO kinematics
Stellar-Based Measurements CO-Based Measurements
Name $D$ $M_{\rm BH}$ $M^{*}/L$ $i$ $M_{\rm BH}$ $M^{*}/L$ $i$ Ref.
(Mpc) ($10^{9}\,M_{\odot}$) [band] ($10^{9}\,M_{\odot}$) [band]
(1) (2) (3) (4) (5) (6) (7) (8) (9)
NGC 315 65.9 $3.0{\pm}0.2$ $2.49{\pm}0.06$ [J] N/A (triaxial) $1.96^{+0.30}_{-0.13}$ $1.86{\pm}0.01$ [J] $74.1^{\circ}{\pm}0.1^{\circ}$ (a,b)
NGC 524 23.3 $0.83^{+0.09}_{-0.04}$ $5.8{\pm}0.4$ [I] $20^{\circ}$ (fixed) $0.40^{+0.12}_{-0.07}$ $5.7{\pm}0.3$ [I] $20^{\circ}$ (fixed) (c,d)
NGC 1332 22.3 $1.45{\pm}0.2$ $7.08{\pm}0.39$ [R] $90^{\circ}$ (fixed) $0.664^{+0.065}_{-0.063}$ $7.83$ [R] $85.2^{\circ}$ (e,f)
NGC 3258 31.9 $2.2{\pm}0.2$ $2.5{\pm}0.1$ [H] $48^{\circ}$ (fixed) $2.249{\pm}0.27$cdots [H] $27.5^{\circ}{-}49.3^{\circ}$ (g,h)
NGC 4697 11.4 $0.18{\pm}0.05$ $4.3{\pm}0.3$ [V] $90^{\circ}$ (fixed) $0.13^{+0.003}_{-0.006}$ $2.14^{+0.04}_{-0.05}$ [i] $76.1^{\circ}{}^{+0.5^{\circ}}_{-0.4^{\circ}}$ (i,j)
NGC 4751 26.9 $1.4{\pm}0.1$ $12.2^{+0.6}_{-0.7}$ [R] $90^{\circ}$ (fixed) $3.43^{+0.16}_{-0.16}$ $2.68{\pm}0.11$ [H] $78.7^{\circ}{}^{+0.1^{\circ}}_{-0.1^{\circ}}$ (k,l)
NGC 6861 27.3 $2.0{\pm}0.2$ $6.1_{-0.1}^{+0.2}$ [I] $90^{\circ}$ (fixed) $1{-}3$ $2.14{-}2.52$ [H] $72.7^{\circ}{-}73.6^{\circ}$ (k,m)
Table 2: Column 1: galaxy name. Column 2: distance. Different values are assumed in the CO vs. stellar studies for NGC 315 (this paper) and NGC 4697; the surface brightness fluctuation distance is adopted here and all measurements are scaled to this value. Column 3: black hole mass from stellar-based measurements. Column 4: stellar mass-to-light ratio for the stellar-based measurements (band indicated in square brackets). Column 5: inclination angle assumed in axisymmetric orbit modeling; only NGC 315 is modeled with a triaxial orbit code. Column 6: black hole mass from CO-based measurements. Column 7: stellar mass-to-light ratio for the CO-based measurements (band indicated in square brackets). Column 8: inclination from CO-based measurements. Column 9: references. (a) This work, (b) Boizelle et al. (2021), (c) Krajnović et al. (2009), (d) Smith et al. (2019), (e) Rusli et al. (2011), (f) Barth et al. (2016), (g) Waters et al. (2024), (h) Boizelle et al. (2019), (i) Schulze & Gebhardt (2011), (j) Davis et al. (2017), (k) Rusli et al. (2013), (l) Dominiak et al. (2024), (m) Kabasares et al. (2022). The reported errors for NGC 524 parameters are $3\sigma$ regions; we divide them by $3$ to approximate the $1\sigma$ regions, as quoted for all other galaxies.
5 Summary
We have performed triaxial stellar orbit modeling of the massive elliptical galaxy NGC 315 using photometric data and ${\sim}2300$ spatially resolved stellar kinematic measurements in 304 bins covering a radial range of ${\sim}0.3^{\prime\prime}$ to $30^{\prime\prime}$ from the MASSIVE survey as constraints. After searching over ${\sim}15,000$ galaxy models with an efficient Bayesian scheme, we are able to simultaneously constrain NGC 315’s $M_{\mathrm{BH}}$, $M^{*}/L$, dark matter halo mass, and intrinsic shape parameters.
We find NGC 315 to be a triaxial and highly prolate galaxy with a triaxiality parameter $T=0.89{\pm}0.02$, hosting a SMBH with $M_{\mathrm{BH}}=(3.0{\pm}0.3){\times}10^{9}M_{\odot}$. At this dynamically inferred mass, the NGC 315 SMBH is located below the mean $M_{\mathrm{BH}}{-}\sigma$ and $M_{\mathrm{BH}}{-}M_{\rm bulge}$ scaling relations formed by other local SMBHs and their host galaxies, but it lies within the intrinsic scatter of these relations. In comparison, the SMBH mass inferred from CO kinematics is $M_{\mathrm{BH}}=(1.96^{+0.30}_{-0.13}){\times}10^{9}M_{\odot}$ (scaled to our distance).
The orbit-based $M_{\mathrm{BH}}$ determination reported in this paper adds a measurement to a small but growing sample of galaxies for which the mass of the central SMBH has been measured using more than one dynamical tracers. Comparing $M_{\mathrm{BH}}$ values for a sample of seven galaxies with both CO-based and stellar-orbit based measurements, we find that the one-to-one relation with an intrinsic scatter term has roughly the same support as a linear relation with free slope, intercept, and intrinsic scatter. The current data therefore do not indicate statistically significant biases between the masses inferred from the two methods.
At our best estimates of $M_{\mathrm{BH}}=3.0{\times}10^{9}M_{\odot}$ and $D=65.9$ Mpc, the NGC 315 SMBH has an angular shadow size of $\theta=2\sqrt{27}GM_{\mathrm{BH}}/c^{2}D\approx 4.7\,\mu{\rm as}$. Together with its relatively high millimeter flux, NGC 315 is a prime candidate target for future event horizon scale imaging missions (e.g., Johnson et al. 2024; Zhang et al. 2024; Ben Zineb et al. 2024). A successful measurement of the shadow size would provide another independent estimate of this SMBH’s mass.
We acknowledge support from NSF-AST-2206219, NSF AST-2206307, and the Heising-Simons Foundation. This work used the Extreme Science and Engineering Discovery Environment (XSEDE) at the San Diego Supercomputing Center through allocation AST180041, which is supported by NSF grant ACI-1548562. Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing. This work is based in part on data obtained at the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation on behalf of the Gemini partnership: the National Science Foundation (United States), the National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). This work is based in part on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. These observations are associated with program GO-14219.
6 Appendix
Figure 7: Comparison of the deprojected 3D luminosity density for the three MGE models of NGC 315 presented in Boizelle et al. (2021). The large bumps in the central 3D luminosity densities of model B2 and B3 are an artifact of the small widths of the central Gaussian component of these two models, $\sigma^{\prime}=0\farcs 178$ and $0\farcs 119$, respectively. In comparison, model A has $\sigma^{\prime}=0\farcs 580$ for the central component and is well behaved upon deprojection.